# Fast-Convergent Proximity Graphs for Approximate Nearest Neighbor Search

### v4

@@ Abstract.
Approximate nearest neighbor (ANN) search in high-dimensional metric spaces is a fundamental problem with many applications. Over the past decade, proximity graph (PG)-based indexes have demonstrated superior empirical performance over alternatives. However, these methods often lack theoretical guarantees regarding the quality of query results, especially in the worst-case scenarios. In this paper, we introduce the $\alpha$-convergent graph ($\alpha$-CG), a new PG structure that employs a carefully designed edge pruning rule. This rule eliminates candidate neighbors for each data point $p$ by applying the shifted-scaled triangle inequalities among $p$, its existing out-neighbors, and new candidates. If the distance between the query point $q$ and its exact nearest neighbor $v^{*}$ is at most $\tau$ for some constant $\tau>0$, our $\alpha$-CG finds the exact nearest neighbor in poly-logarithmic time, assuming bounded intrinsic dimensionality for the dataset; otherwise, it can find an ANN in the same time. To enhance scalability, we develop the $\alpha$-convergent neighborhood graph ($\alpha$-CNG), a practical variant that applies the pruning rule locally within each point’s neighbors. We also introduce optimizations to reduce the index construction time. Experimental results show that our $\alpha$-CNG outperforms existing PGs on real-world datasets. For most datasets, $\alpha$-CNG can reduce the number of distance computations and search steps by over 15% and 45%, respectively, when compared with the best-performing baseline.
@@
Figure 1. The edge pruning rules of existing MRNG, $\tau$-MG, Vamana, and our $\alpha$-CG. Given a data point $p$, a candidate $u$ is pruned if $p$ already has an out-neighbor $v$ falling within the intersection of $B(p,\mathit{\delta}(p,u))$ and $B(u,r)$.
@@
Method Edge pruning rule $\delta(q,v^{*})=0$ $\delta(q,v^{*})<\tau$ $\delta(q,v^{*})>\tau$ Routing Length Average Time Complexity Worst-case Time Complexity
MRNG (fxwc19) Triangle Inequality
@@
Vamana (sds+19) scaled Tri. Ine.
@@
$\alpha$-CG (Ours) shifted-scaled Tri. Ine.
$O(\log_{\alpha}\Delta)^{\dagger}$ $O\left((\alpha\cdot\tau)^{d}\log\Delta\log_{\alpha}\Delta\right)^{\dagger}$
@@
1. Introduction
Approximate nearest neighbor search is a fundamental component in various applications such as image retrieval (kg09; pci+07), recommendation systems (ddgr07; bnmn16), and data mining (hff+17; i16). Let $P$ be a set of $n$ data points from a metric space $(\mathcal{X},\mathit{\delta})$ where $\mathcal{X}$ is a set of points and $\mathit{\delta}:\mathcal{X}\times\mathcal{X}\rightarrow_{\geq 0}$ is a distance function. We want to construct a data structure on $P$ that supports the following nearest neighbor (NN) query: given a query point $q\in\mathcal{X}$, return the point $v^{*}\in P$ closest to $q$. Due to the “curse of dimensionality”, all known space-efficient solutions for NN search have search times that grow exponentially with the dimensionality of the metric space. To enhance query efficiency, researchers often resort to $c$-approximate nearest neighbor (ANN) queries for some constant $c>1$, which balance query accuracy and search time by returning a point whose distance to $q$ is within a $c$-factor of the exact NN.
In recent years, numerous studies (my20; fxwc19; pcc+23; sds+19; wxyw21; aep25) have demonstrated that proximity graph-based solutions exhibit superior empirical performance on real-world data over other indexes (e.g., trees (wwj+13; laja08) and inverted index (bl14; jds10)). A proximity graph (PG) is a simple directed graph, where each vertex represents a data point $p\in P$ and edges are connected based on particular geometric properties (usually close neighbors in distance). Given a query point $q$, a simple greedy routing is performed to find an ANN of $q$: starting from a fixed or random entry vertex, at each step, it explores the nearest out-neighbor of the currently visited vertex to $q$. This process traverses a search path until no closer nodes can be identified, returning the closest visited point as the answer [^1] We can generalize the algorithm to find $k$ ANNs by maintaining a sorted list, known as the beam search algorithm (see Section 2.2).. The running time of the algorithm is bounded by the total out-degree of the nodes on the search path. Therefore, to ensure efficient query performance, the goal is to construct a sparse graph (with a low maximum out-degree) that allows the search algorithm to converge to the ANNs of $q$ quickly.
@@ 1.1. Existing PGs and Their Limitations
Most methods for constructing PGs follow a common framework. For each data point $p$, we first obtain a candidate set $\mathcal{V}$ of points from $P$ that are close to $p$, through greedy search on a base graph (e.g., an approximate $K$-NN graph (fxwc19; pcc+23)) or an intermediate graph over a subset of $P$ (my20; sds+19). A small subset of $\mathcal{V}$ is then selected as the out-neighbors of $p$ to maintain graph connectivity. Crucially, the resulting graph should have the “shortcutable” property: for any query point $q$, if data point $p$ is not an ANN of $q$, $p$ should have an out-neighbor that is closer to $q$ than $p$. To achieve this, existing PG solutions introduce edge pruning rules to eliminate unnecessary candidates in $\mathcal{V}$ in the following procedure. For each point $u\in\mathcal{V}$ processed in ascending order of the distances to $p$:
@@ 1.2. Our Contributions
    A new proximity graph with enhanced guarantees. We propose a new PG, the $\alpha$-convergent graph ($\alpha$-CG), in which we set $\mathcal{V}=P\setminus\{p\}$ for each $p\in P$, and then utilize the shifted-scaled triangle inequality that incorporates both parameters $\tau$ and $\alpha$ to prune unnecessary candidates. Specifically, as illustrated in Figure 1d, we set $r=\frac{1}{\alpha}(\mathit{\delta}(p,u)-(\alpha+1)\tau)$, which reduces the radius of the ball $B(u,\mathit{\delta}(p,u))$ by first subtracting $(\alpha+1)\tau$ from $\mathit{\delta}(p,u)$ and then scaling a multiplicative factor of $\alpha$. The intersection area is strictly smaller than that of the MRNG and Vamana.
    With this new edge pruning rule, we prove that the $\alpha$-CG admits non-trivial theoretical guarantees. When $\mathit{\delta}(q,v^{*})\leq\tau$, our $\alpha$-CG can find the exact NN of $q$ in $O((\alpha\tau)^{d}\log\Delta\log_{\alpha}\Delta)$ time, thereby achieving the first algorithm that can find the exact NN in poly-logarithmic time when the doubling dimension is constant (see Table 1). A critical property of our PG is that after each hop, the distance from the next visited node to $q$ is reduced by an $\alpha$-multiplicative factor, and the search algorithm terminates in $O(\log_{\alpha}\Delta)$ steps. While $\tau$-MG and the slow preprocessing version of Vamana terminates in $O(n)$ and $O(\log_{\alpha}\frac{\Delta}{(\alpha-1)\epsilon})$ steps, respectively. When $\mathit{\delta}(q,v^{*})>\tau$, a greedy routing on $\alpha$-CG can find an $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN of $q$ for any $\epsilon>0$, with the same query accuracy as that of Vamana. The space, query time, and construction time of $\alpha$-CG match those of Vamana (ix23), up to an $O(\tau^{d})$-factor (see Theorem 5).
@@
    Empirical analysis shows that the parameter $\alpha$ greatly affects the performance of our graph. Increasing $\alpha$ reduces the number of search steps but results in higher out-degrees in the PG. This trade-off complicates the selection of an optimal $\alpha$, a problem that subsequent works (sds+19; ssks21; gks+23; jkgsa22) on Vamana have overlooked. We propose an adaptive local pruning strategy that adjusts $\alpha$ for each node locally during graph construction. Starting from a small value, we gradually increase $\alpha$ and prune candidates until the node’s out-degree reaches a predefined threshold, preserving long-distance shortcut edges and maintaining graph connectivity. To construct our graph efficiently, we implement a distance-reusing mechanism that stores and reuses intermediate computation results to accelerate the adaptive pruning process, along with a lazy pruning strategy that significantly reduces the number of pruning operations.
@@ 2. Preliminaries
    This section will introduce the (approximate) nearest neighbor search problem and basic notations. Then, we provide an overview of proximity graphs and the concept of doubling dimension.
@@ 2.1. Problem Setting and Basic Notations
    A metric space $(\mathcal{X},\mathit{\delta})$ consists of a set $\mathcal{X}$ of points and a distance function $\mathit{\delta}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ satisfying the inequality $\mathit{\delta}(p_{1},p_{2})+\mathit{\delta}(p_{2},p_{3})\geq\mathit{\delta}(p_{1},p_{3}),\>\forall p_{1},p_{2},p_{3}\in\mathcal{X}$. Let $P\subseteq\mathcal{X}$ be a set of $n$ data points from $\mathcal{X}$. Given a query point $q$, a point $v^{*}\in P$ is a nearest neighbor (NN) of $q$ if $\mathit{\delta}(v^{*},q)\leq\mathit{\delta}(p,q)$ for all $p\in P$; while a point $p\in P$ is a $c$-approximate nearest neighbor ($c$-ANN) of $q$ for some constant $c>1$ if $\mathit{\delta}(p,q)\leq c\cdot\mathit{\delta}(p^{\prime},q)$ for all $p^{\prime}\in P$. We want to preprocess $P$ into a data structure with a small space that can answer exact NN or ANN queries efficiently.
@@
    Following the convention of previous works (ix23; fxwc19; pcc+23), this paper will assume that each distance computation consumes constant time. The query time of the beam search algorithm is asymptotically bounded by the total number of distance computations, which is the total out-degree of the visited hop vertices.
@@ 2.3. Doubling Dimension
    Given a point $p$ and a real value $r>0$, we will use $B(p,r)$ to represent the set containing all the points in $\mathcal{X}$ whose distance to $p$ is at most $r$; we refer to $B(p,r)$ as a ball centered at $p$ with radius $r$. To analyze the performance of PG-based algorithms, we introduce the notion of doubling dimension, which is often used to measure the “intrinsic dimensionality” of high-dimensional point sets (gkl03; bkl06; kl04):
@@
    The doubling dimension generalizes the Euclidean dimension. For any set $D\subseteq^{m}$, the doubling dimension of $D$ is O(m)O(m). [^4] https://en.wikipedia.org/wiki/Doubling_space For example, when $m=2$, any ball of radius $r$ can be covered by 7 balls with radius $r/2$, meaning that the doubling constant of $D$ is at most $\log_{2}7$. Moreover, empirical studies (fk94; ix23) showed that the doubling dimension of real data sets is often smaller than their ambient dimension, e.g., dimension $d$ of the Euclidean space. Following the previous studies on doubling dimension (cg06; hm06; kl04), this paper assumes that the doubling dimension of input $P$ — denoted as $d$ — is $O(1)$, namely, $P$ has a low doubling dimension. Based on the definition of doubling dimension, we can derive (ix23):
@@
    3. Alpha-Convergent Graph
    This section presents a new PG that offers stronger theoretical guarantees on query time and accuracy. Similar to Vamana and $\tau$-MG, its construction time is $\Omega(n^{2})$. We will provide a practical variant in Section 4 to reduce the construction time.
    3.1. A New Pruning Strategy
    Definition 0 (Neighbor Selection (aep25)).
    Definition 0 (Shifted-Scaled Triangle Inequality).
    where $\alpha>1$ is a parameter.
@@ 3.2. The $\alpha$-Convergent Graph
    Based on our new edge pruning rule, we define the $\alpha$-convergent graph $G$ ($\alpha$-CG) of $P$ as follows:
@@
    Figure 2. Illustration of $\alpha$-reducible property
@@
    Combining the above with the inequality $\mathit{\delta}(p,q)\geq\mathit{\delta}(p,v^{*})-\mathit{\delta}(q,v^{*})$ (by definition of the metric space), we have
@@
    where the last inequality used the assumption that $\mathit{\delta}(q,v^{*})\leq\tau$. On the other hand, we have (by definition of the metric space)
@@
    See an illustration of the property in Figure 2. Given any query $q$ satisfying $\mathit{\delta}(q,v^{*})\leq\tau$, Lemma 4 implies that greedy routing from any starting vertex quickly converges to the exact NN of $q$: at each step, if the current vertex is not $v^{*}$, then either the next hop reaches $v^{*}$ or the distance from the next vertex to $q$ decreases by an $\alpha$ factor.

### v5

@@ Abstract.
Approximate nearest neighbor (ANN) search in high-dimensional metric spaces is a fundamental problem with many applications. Over the past decade, proximity graph (PG)-based indexes have demonstrated superior empirical performance over alternatives. However, these methods often lack theoretical guarantees regarding the quality of query results, especially in the worst-case scenarios. In this paper, we introduce the $\alpha$-convergent graph ($\alpha$-CG), a new PG structure that employs a new carefully designed edge pruning rule. If the distance between the query point $q$ and its exact nearest neighbor $v^{*}$ is at most $\tau$ for some constant $\tau>0$, our $\alpha$-CG finds the exact nearest neighbor in poly-logarithmic time, assuming bounded intrinsic dimensionality for the dataset; otherwise, it can find an ANN in the same time. To enhance scalability, we develop the $\alpha$-convergent neighborhood graph ($\alpha$-CNG), a practical variant that applies the pruning rule locally within each point’s neighbors. We also introduce optimizations to reduce the index construction time. Experimental results show that our $\alpha$-CNG outperforms existing PGs on real-world datasets. For most datasets, $\alpha$-CNG can reduce the number of distance computations and search steps by over 15% and 45%, respectively, when compared with the best-performing baseline.
Proximity Graphs, Approximate Nearest Neighbor Search, Data Structures
journal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8copyright: ccjournal: PACMMODjournalyear: 2026journalvolume: 4journalnumber: 1 (SIGMOD)article: 36publicationmonth: 2doi: 10.1145/3786650ccs: Theory of computation Data structures design and analysisccs: Information systems Information retrieval query processing
@@
Figure 1. The edge pruning rules of existing MRNG, $\tau$-MG, Vamana, and $\alpha$-CG. Given a data point $p$, a candidate $u$ is pruned if $p$ already has an out-neighbor $v$ falling within the intersection of $B(p,\mathit{\delta}(p,u))$ and $B(u,r)$.
1. Introduction
Approximate nearest neighbor search is a fundamental component in various applications such as image retrieval (Kulis and Grauman, 2009; Philbin et al., 2007), recommendation systems (Das et al., 2007; Boytsov et al., 2016), and data mining (Huang et al., 2017; Iwasaki, 2016). Let $P$ be a set of $n$ data points from a metric space $(\mathcal{X},\mathit{\delta})$ where $\mathcal{X}$ is a set of points and $\mathit{\delta}:\mathcal{X}\times\mathcal{X}\rightarrow_{\geq 0}$ is a distance function. We want to construct a data structure on $P$ that supports the following nearest neighbor (NN) query: given a query point $q\in\mathcal{X}$, return the point $v^{*}\in P$ closest to $q$. Due to the “curse of dimensionality”, all known space-efficient solutions for NN search have search times that grow exponentially with the dimensionality of the metric space. To enhance query efficiency, researchers often resort to $c$-approximate nearest neighbor (ANN) queries for some constant $c>1$, which balance query accuracy and search time by returning a point whose distance to $q$ is within a $c$-factor of the exact NN.
In recent years, numerous studies (Malkov and Yashunin, 2020; Fu et al., 2019; Peng et al., 2023; Subramanya et al., 2019; Wang et al., 2021; Azizi et al., 2025) have demonstrated that proximity graph-based solutions exhibit superior empirical performance on real-world data over other indexes (e.g., trees (Wang et al., 2013; Lejsek et al., 2008) and inverted index (Babenko and Lempitsky, 2014; Jegou et al., 2010)). A proximity graph (PG) is a simple directed graph, where each vertex represents a data point $p\in P$ and edges are connected based on particular geometric properties (usually close neighbors in distance). Given a query point $q$, a simple greedy routing is performed to find an ANN of $q$: starting from a fixed or random entry vertex, at each step, it explores the nearest out-neighbor of the currently visited vertex to $q$. This process traverses a search path until no closer nodes can be identified, returning the closest visited point as the answer [^1] We can generalize the algorithm to find $k$ ANNs by maintaining a sorted list, known as the beam search algorithm (see Section 2.2).. The running time of the algorithm is bounded by the total out-degree of the nodes on the search path. Therefore, to ensure efficient query performance, the goal is to construct a sparse graph (with a low maximum out-degree) that allows the search algorithm to converge to the ANNs of $q$ quickly.
@@ 1. Introduction
Method Pruning radius $\delta(q,v^{*})=0$ $\delta(q,v^{*})<\tau$ $\delta(q,v^{*})>\tau$ Routing Length Avg. Time Worst-case Time
MRNG (Fu et al., 2019) $d(p,u)$
@@
Vamana (Subramanya et al., 2019) $\frac{1}{\alpha}d(p,u)$
@@
$\alpha$-CG (Ours) $\frac{1}{\alpha}d(p,u)-(\alpha+1)\tau$
$O(\log_{\alpha}\Delta)^{\dagger}$ $O\!\left((\alpha\tau)^{d}\log\Delta\log_{\alpha}\Delta\right)^{\dagger}$
@@ 1.1. Existing PGs and Their Limitations
Most methods for constructing PGs follow a common framework. For each data point $p$, we first obtain a candidate set $\mathcal{V}$ of points from $P$ that are close to $p$, through greedy search on a base graph (e.g., an approximate $K$-NN graph (Fu et al., 2019; Peng et al., 2023)) or an intermediate graph over a subset of $P$ (Malkov and Yashunin, 2020; Subramanya et al., 2019). A small subset of $\mathcal{V}$ is then selected as the out-neighbors of $p$ to maintain graph connectivity. Crucially, the resulting graph should have the “shortcutable” property: for any query point $q$, if data point $p$ is not an ANN of $q$, $p$ should have an out-neighbor that is closer to $q$ than $p$. To achieve this, existing PG solutions introduce edge pruning rules to eliminate unnecessary candidates in $\mathcal{V}$ in the following procedure. For each point $u\in\mathcal{V}$ processed in ascending order of the distances to $p$:
@@ 1.2. Our Contributions
    A new proximity graph with enhanced guarantees. We propose a new PG, the $\alpha$-convergent graph ($\alpha$-CG), in which we set $\mathcal{V}=P\setminus\{p\}$ for each $p\in P$, and then utilize a carefully designed edge pruning rule that incorporates both parameters $\tau$ and $\alpha$ to eliminate unnecessary candidates. Specifically, as illustrated in Figure 1d, we set $r=\frac{1}{\alpha}(\mathit{\delta}(p,u)-(\alpha+1)\tau)$, which reduces the radius of the ball $B(u,\mathit{\delta}(p,u))$ by first subtracting $(\alpha+1)\tau$ from $\mathit{\delta}(p,u)$ and then scaling a multiplicative factor of $\alpha$. The intersection area is strictly smaller than that of the MRNG and Vamana.
    With this new edge pruning rule, we prove that the $\alpha$-CG admits non-trivial theoretical guarantees. When $\mathit{\delta}(q,v^{*})\leq\tau$, our $\alpha$-CG can find the exact NN of $q$ in $O((\alpha\tau)^{d}\log\Delta\log_{\alpha}\Delta)$ time, thereby achieving the first algorithm that can find the exact NN in poly-logarithmic time when the doubling dimension is constant (see Table 1). A critical property of our PG is that after each hop, the distance from the next visited node to $q$ is reduced by an $\alpha$-multiplicative factor, and the search algorithm terminates in $O(\log_{\alpha}\Delta)$ steps. While $\tau$-MG and the slow preprocessing version of Vamana terminates in $O(n)$ and $O(\log_{\alpha}\frac{\Delta}{(\alpha-1)\epsilon})$ steps, respectively. When $\mathit{\delta}(q,v^{*})>\tau$, a greedy routing on $\alpha$-CG can find an $(\frac{\alpha+1}{\alpha-1}+\epsilon)$-ANN of $q$ for any $\epsilon>0$, with the same query accuracy as that of Vamana. The space, query time, and construction time of $\alpha$-CG match those of Vamana (Indyk and Xu, 2023), up to an $O(\tau^{d})$-factor (see Theorem 5).
@@
    Empirical analysis shows that the parameter $\alpha$ greatly affects the performance of our graph. Increasing $\alpha$ reduces the number of search steps but results in higher out-degrees in the PG. This trade-off complicates the selection of an optimal $\alpha$, a problem that subsequent works (Subramanya et al., 2019; Singh et al., 2021; Gollapudi et al., 2023; Jaiswal et al., 2022) on Vamana have overlooked. We propose an adaptive local pruning strategy that adjusts $\alpha$ for each node locally during graph construction. Starting from a small value, we gradually increase $\alpha$ and prune candidates until the node’s out-degree reaches a predefined threshold, preserving long-distance shortcut edges and maintaining graph connectivity. To construct our graph efficiently, we implement a distance-reusing mechanism that stores and reuses intermediate computation results to accelerate the adaptive pruning process, along with a lazy pruning strategy that significantly reduces the number of pruning operations.
@@ 2.1. Problem Setting and Basic Notations
    A metric space $(\mathcal{X},\mathit{\delta})$ consists of a set $\mathcal{X}$ of points and a distance function $\mathit{\delta}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$ satisfying the triangle inequality $\mathit{\delta}(p_{1},p_{2})+\mathit{\delta}(p_{2},p_{3})\geq\mathit{\delta}(p_{1},p_{3}),\>\forall p_{1},p_{2},p_{3}\in\mathcal{X}$. Let $P\subseteq\mathcal{X}$ be a set of $n$ data points from $\mathcal{X}$. Given a query point $q$, a point $v^{*}\in P$ is a nearest neighbor (NN) of $q$ if $\mathit{\delta}(v^{*},q)\leq\mathit{\delta}(p,q)$ for all $p\in P$; while a point $p^{\prime}\in P$ is a $c$-approximate nearest neighbor ($c$-ANN) of $q$ for some constant $c>1$ if $\mathit{\delta}(p^{\prime},q)\leq c\cdot\mathit{\delta}(p,q)$ for all $p\in P$. We want to preprocess $P$ into a data structure with a small space that can answer exact NN or ANN queries efficiently.
@@
    Following the convention of previous works (Indyk and Xu, 2023; Fu et al., 2019; Peng et al., 2023), this paper will assume that each distance computation consumes constant time. The query time of the beam search algorithm is asymptotically bounded by the total number of distance computations, which is the total out-degree of the visited hop vertices.
@@ 2.3. Doubling Dimension
    Given a point $p$ and a real value $r>0$, we will use $B(p,r)$ to represent the set containing all the points in $\mathcal{X}$ whose distance to $p$ is at most $r$; we refer to $B(p,r)$ as a ball centered at $p$ with radius $r$. To analyze the performance of PG-based algorithms, we introduce the notion of doubling dimension, which is often used to measure the “intrinsic dimensionality” of high-dimensional point sets (Gupta et al., 2003; Beygelzimer et al., 2006; Krauthgamer and Lee, 2004):
@@
    The doubling dimension generalizes the Euclidean dimension. For any set $D\subseteq^{m}$, the doubling dimension of $D$ is O(m)O(m). [^4] https://en.wikipedia.org/wiki/Doubling_space For example, when $m=2$, any ball of radius $r$ can be covered by 7 balls with radius $r/2$, meaning that the doubling constant of $D$ is at most $\log_{2}7$. Moreover, empirical studies (Faloutsos and Kamel, 1994; Indyk and Xu, 2023) showed that the doubling dimension of real data sets is often smaller than their ambient dimension, e.g., dimension $d$ of the Euclidean space. Following the previous studies on doubling dimension (Cole and Gottlieb, 2006; Har-Peled and Mendel, 2006; Krauthgamer and Lee, 2004), this paper assumes that the doubling dimension of input $P$ — denoted as $d$ — is $O(1)$, namely, $P$ has a low doubling dimension. Based on the definition of doubling dimension, we can derive (Indyk and Xu, 2023):
@@
    3. Alpha-Convergent Graph
    This section presents a new PG, called the $\alpha$-convergent graph ($\alpha$-CG), which offers stronger theoretical guarantees on query time and accuracy. Similar to Vamana and $\tau$-MG, its construction time is $\Omega(n^{2})$. We will provide a practical variant in Section 4 to reduce the construction time.
    3.1. A New Pruning Strategy
    Definition 0 (Neighbor Selection (Azizi et al., 2025)).
    Definition 0 (Edge pruning rule of $\alpha$-CG).
    where $\alpha>1$ is a parameter.
@@ 3.2. The $\alpha$-Convergent Graph
    Based on our new edge pruning rule, we define the $\alpha$-CG $G$ of $P$ as follows:
@@
    (a) $\alpha$-reducible property
    (b) Out-degree analysis: Case 1
    (c) Out-degree analysis: Case 2
    Figure 2. Geometric illustrations of the pruning rule. (a) Illustration of the $\alpha$-reducible property. (b,c) Two cases in the out-degree analysis.
@@
    Combining the above with the triangle inequality $\mathit{\delta}(p,q)\geq\mathit{\delta}(p,v^{*})-\mathit{\delta}(q,v^{*})$, we have
@@
    where the last inequality used the assumption that $\mathit{\delta}(q,v^{*})\leq\tau$. On the other hand, according to the triangle inequality, we have
@@
    See an illustration of the property in Figure 2(a). Given any query $q$ satisfying $\mathit{\delta}(q,v^{*})\leq\tau$, Lemma 4 implies that greedy routing from any starting vertex quickly converges to the exact NN of $q$: at each step, if the current vertex is not $v^{*}$, then either the next hop reaches $v^{*}$ or the distance from the next vertex to $q$ decreases by an $\alpha$ factor.